complete the square, and find the roots of the quadratic equation.
x^2 - 2/3x = 0

Answer by jcmtnez(53)

You don't need to complete the square since x^2-2/3x=0 can be written as follow:
x(x-2/3)=0 Therefore x=0
And x-2/3=0, x=2/3 which is the second solution.